Determine the amount of theoretical oxygen and theoretical air needed to burn
a.) 60 kg of butane (C4H10)
b.) 20 kg of propane (C5H12)
c.) 80 kg of octane (C8H18)
1
Expert's answer
2020-12-06T13:52:48-0500

A. 2C4H10 + 13O2 ----> 8CO2 + 10H2O

Molar mass of C4H10 = (12x4)+10= 58gmol-1

molar mass of O2= 16x2=32gmol-1

2(58)g of C4H10 required 13(32)g of Oxygen

Therefore 60Kg of C4H10 will require 416x60/116 = 215.2Kg of Oxygen

Since oxygen is 21% of Air

Theoretical amount of air needed= 100x 215.2/21 = 1025Kg of air


B. 2C5H12 + 17O2 -----> 12CO2 + 10H2O

Molar mass of C5H12 = (12x5) + 12= 72gmol-1

2(72)g of C5H12 required 17(32)g of Oxygen

Therefore, 20Kg of C5H12will require 544x20/144=75.6Kg of oxygen gas

Amount of air= 100x75.6/21= 360Kg of air


C. 2C8H18 + 25O2 -----> 16CO2 + 18H2O

Molar mass of C8H18 = (12x8) + 18 = 114gmol-1

2(114)g of C8H18 requires 25(32)g of Oxygen gas

80Kg of C8H18 will require 800x80/288 = 280.7Kg of Oxygen gas

Amount of air needed= 100x280.7/21= 1337Kg of air.